Air United, Inc. manufactures two products: missile range instruments and space pressure gauges. During April, 54 range instruments and 303 pressure gauges were produced, and overhead costs of $82,400 were estimated. An analysis of estimated overhead costs reveals the following activities.

Activities

Cost Drivers

Total Cost

1. Materials handling Number of requisitions

$32,960

2. Machine setups Number of setups

26,780

3. Quality inspections Number of inspections

22,660

$82,400


The cost driver volume for each product was as follows.

Cost Drivers

Instruments

Gauges

Total

Number of requisitions 425 605 1,030
Number of setups 220 295 515
Number of inspections 235 280 515

Determine the overhead rate for each activity.

Overhead Rate

Materials handling

$enter a dollar amount

Machine setups

$enter a dollar amount

Quality inspections

$enter a dollar amount

eTextbook and Media

  

  

Assign the manufacturing overhead costs for April to the two products using activity-based costing. (Round per unit answers to 2 decimal places, e.g. 12.25.)

Instruments

Gauges

Total cost assigned

$enter a dollar amount

$enter a dollar amount

Overhead cost per Unit

$enter a dollar amount rounded to 2 decimal places

$enter a dollar amount rounded to 2 decimal places
Answer #1

1. Overhead Rate for Each Activity

Overhead Overhead Cost ÷ Cost Driver Volume = Overhead Rate
Material Handling $32,960 ÷ 1,030 = $32
Machine Setups $26,780 ÷ 515 = $52
Quality Inspection $22,660 ÷ 515 = $44
Total $82,400

For Student understanding ( How to interpret)

  • Material Handling overhead rate is $32 per requisitions
  • Machine Setups overhead rate is $52 per Setups
  • Quality Inspection Overhead rate is $44 per inspections

2. Assignment of manufacturing overhead costs to two product using Activity based costing

Particulars Instruments Gauges
Total Cost Assigned $35,380.00 $47,020.00
Overhead Cost Per Unit $ 655.19 $ 155.18

Working

(i) Cost assigned to Instruments

Activity name Activity Rates Activity ABC Cost
(A) (B) (A x B)
Material handling cost $32.00                      425.00 $13,600.00
Printing $52.00                      220.00 $11,440.00
Product design $44.00                      235.00 $10,340.00
Total Overheads assigned $35,380.00
Production                   54.00
Overhead cost per unit $ 655.19

(ii) Cost assigned to Gauges

Activity name Activity Rates Activity ABC Cost
(A) (B) (A x B)
Material handling cost $32.00                      605.00 $19,360.00
Printing $52.00                      295.00 $15,340.00
Product design $44.00                      280.00 $12,320.00
Total Overheads assigned $47,020.00
Production                   303.00
Overhead cost per unit $ 155.18

*Highlighted portion are main Answers of Question

.